OpenStudy (inexilewetrust):

I am not sure how to do this, can someone explain to me how to solve this? A single 1,300 kg cargo train car is rolling along a train track at 2.0 m/s when 400 kg of coal is dropped vertically into it. What is its velocity right afterward? Assume a closed system. 1.5 m/s 2.0 m/s 6.5 m/s 0.0011 m/s

OpenStudy (vincent-lyon.fr):

Use conservation of momentum just before and after the collision.

OpenStudy (inexilewetrust):

How exactly do i do that?

OpenStudy (shamim):

U hv to use the formula m1u1+m2u2=m1v1+m2v2

OpenStudy (shamim):

Response plz for better help

OpenStudy (shamim):

Given m1=1300kg U1=2m/s

OpenStudy (shamim):

m2=1300+400=1700kg u2=0 v1=v2=?

OpenStudy (vincent-lyon.fr):

You can also write : m1u1+m2u2=(m1+m2)v

OpenStudy (inexilewetrust):

Thanks guys! I finished the test way before anybody responded. The formulas did help and I really appreciate it! I will write it down.